OpenStudy (anonymous):

2.The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the acceleration due to gravity, and t = time. Solve this equation for g, and use your formula to determine the acceleration due to gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (anonymous):

help please

OpenStudy (cwrw238):

first divide both sides by 0.5t^2 this will give you g in terms of d and t then plug in the given values of d and t into this formula and work out g

OpenStudy (anonymous):

so it will look likeeeeeeee \[d/0.5t ^{2}?\]

OpenStudy (cwrw238):

g = d / (0.5t^2) g = 2d / t^2 t = 10 when d = 490 so g = 2*490 / 10^2

OpenStudy (anonymous):

so where do you get the 2 fro 2*490 from/ @cwrw

OpenStudy (cwrw238):

1 / 0.5 = 2 it looks tidier with 2 in the numerator

OpenStudy (anonymous):

ohhhhhhhh okay thank you so much !!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

can I ask you another question? @cwrw238